OpenStudy (anonymous):

Bob is driving along a straight and level road straight toward a mountain. At some point on his trip he measures the angle of elevation to the top of the mountain and finds it to be 21°44'. Find the height of the mountain to the nearest foot if Bob is 13,474.3 feet from the center of the mountain at the base.

OpenStudy (anonymous):

tan(21+44/60)=-2.63668....?

OpenStudy (anonymous):

No. Everything I do ends up giving me a completely different answer than any of the possible choices. I see what you mean about finding AB I think. It is basically what value of ab divided by 13474.3 gives you tan21′44′′ right?

OpenStudy (anonymous):

the choices are 5453 ft 5353 ft 53,534 ft 535,342 ft I think it is 5353ft.

OpenStudy (anonymous):

:) Haha its ok. You're helping me out a ton. Take your time.

OpenStudy (anonymous):

I keep getting 6871.0846. I did (sin21′44′′)*13474.3 and then doubled it.

OpenStudy (anonymous):

:( That's not one of the choices. The choices are 5453 ft 5353 ft 53,534 ft 535,342 ft Don't ask me how my teacher got those but those are the multiple choice answers he gave me.

OpenStudy (anonymous):

I'm glad I'm not the only one making mistakes. It makes me feel better that you are too. :P

OpenStudy (anonymous):

Now I'm seriously lost. Do I have to multiply by 4 now?

OpenStudy (anonymous):

I have no idea what to do here.

OpenStudy (anonymous):

in triangle BDC , tan x = BD/BC

OpenStudy (anonymous):

in triangle BDC , tan x = BD/BC

OpenStudy (anonymous):

now BD=AD so we can write , tan x= AD/BC -------------> 1.) in triangle ABC, tan21'44''=AB/BC now, AB=2AD so, tan21'44'' = 2AD/BC or , tan21'44'' / 2 = AD/BC -----------------> 2.) since the RHS of 1.) and 2.) are equal, their LHS must be equal, so, tan x = tan21'44''/2 find x. understood uptil now ??

OpenStudy (anonymous):

Wow, I ended up getting 7.1216. I am so completely lost now. What happened to 13474.3?

OpenStudy (anonymous):

is it positive value of negative of x?

OpenStudy (anonymous):

I don't know. I guess negative?

OpenStudy (anonymous):

now , use sin x.

OpenStudy (anonymous):

I don't even know what x is anymore.

OpenStudy (anonymous):

7.1216.

OpenStudy (anonymous):

You mean that was right?

OpenStudy (anonymous):

i have doubt how it can be negative.

OpenStudy (anonymous):

OK well sin(7.1216)=.743584.........

OpenStudy (anonymous):

check your sin value

OpenStudy (anonymous):

I thought you said it was 7.1216?

OpenStudy (anonymous):

sin(7.1216)=0.123976............

OpenStudy (anonymous):

How did you get that? My calculator keeps saying .743584... so on. Maybe that's my problem. My calculator is broken?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=sin%287.1216%29

OpenStudy (anonymous):

OH degree! Oh. I see.

OpenStudy (lgbasallote):

i get this strange feeling that something in that triangle is wrong....

OpenStudy (anonymous):

Oh my f******. I'm never going to get this. I really appreciate your help but at this rate, there is no way I'm picking this up. I am too lost and just completely burnt out on this entire question. I have a million more just like it and its been like 3 hours on the same question. I have no idea what I'm doing with this.

hartnn (hartnn):

i am getting it as 5371.1 tan 21.44 * 13474.3. and nearest option is 5353 ft

OpenStudy (anonymous):

Doing it that way I got 5291.3576. Close enough. Thats the answer I choose 5353. I give up. Thank you guys for your help. I really do appreciate it. I know im not an easy one teach.

hartnn (hartnn):

yeah, don't waste much time on one problem, go to next